An eyewitness to a crime said the license plate to the getaway care began with the letter R, and N, but he couldn't remember the order. The rest of the plate had the numbers 5, 1, and 2, but again couldn't recall in what order. How many license plates must the police investigate to find the car?
Answer by Theo(13342)

there are 2 choices for the first position.
there are 1 choice for the second position.
there are 3 choices for the third position.
there are 2 choices for the fourth position.
there are 1 choice for the fifth position.

the total possibilities are 2 * 1 * 3 * 2 * 1 = 12

those possibilities are:

NR125
NR152
NR215
NR251
NR512
NR521

RN125
RN152
RN215
RN251
RN512
RN521

breaking it down, there are 2 letters that can be in the first two positions in any order, and there are 3 numbers that can be in the last 3 positions in any order.

that gets you 2! * 3! = 2 * 1 * 3 * 2 * 1 = 12 possible arrangements.